OpenStudy (shaleiah):

The speed of sound is 344 m/s. You have built a really fantastic car that can really go fast. If the car can accelerate at 22.4 m/s, how much time till you reach the speed of sound? How many kilometers will you travel before you reach that speed?

OpenStudy (mathstudent55):

Are you sure of the acceleration rate? The unites are incorrect.

OpenStudy (sshayer):

v=u+at u=0 a=22.4 m/s v=344 m/s find t

OpenStudy (shaleiah):

15.35

OpenStudy (sshayer):

\[v^2-u^2=2as\] find s

OpenStudy (mathstudent55):

speed = distance/time acceleration = speed/time

OpenStudy (sshayer):

correction a=22.4 m/s^2

OpenStudy (mathstudent55):

You have the acceleration and the speed, solve for time using the acceleration equation.

OpenStudy (mathstudent55):

Your answer needs to have units, not just a number.

OpenStudy (shaleiah):

do you agree? 15.35 km

OpenStudy (shaleiah):

@sshayer

OpenStudy (sshayer):

15.36 seconds

OpenStudy (sshayer):

it is rounded to two decimals.

OpenStudy (sshayer):

it is time ,so time units.

OpenStudy (sshayer):

for finding second part use formula given by me. distance will be in meters.Change it in kilometers by dividing it by 1000

OpenStudy (sshayer):

i am leaving now.